OpenStudy (anonymous):

derrivative of sec(3x-2)

hartnn (hartnn):

do you know chain rule ?

OpenStudy (anonymous):

no

hartnn (hartnn):

Chain rule : \((f(g(x))' = f'(g(x)) \times g'(x)\) here, first tell me whats the derivative of sec x ?

OpenStudy (anonymous):

sec x times tan x

hartnn (hartnn):

and derivative of 3x-2 ?

OpenStudy (anonymous):

3

hartnn (hartnn):

yes, correct. so, (sec (3x-2))' = sec (3x-2) tan (3x-2) * (3x-2)' =3sec (3x-2) tan (3x-2) got this ?

OpenStudy (anonymous):

yes

hartnn (hartnn):

good! thats your final answer as it cannot be further simplified. Ask for doubts, if any.